OpenStudy (anonymous):

Four integers j, j, k, and n have an average of 0. Which of the following equations MUST be true? k=n k=-j k+n=-2j k+n=0 k+n=j

OpenStudy (anonymous):

3rd

OpenStudy (anonymous):

k+n=-2j? can u please give an explanation of how u got that?

OpenStudy (lgbasallote):

this situation is only possible if J + j + k + n = 0 if i simplify that i have 2j + k + n = 0 if i subtract both sides by 2j k + n = -2j does that help?

OpenStudy (anonymous):

yes